1. Consider a time-homogeneous Markov chain X)n, such that P= 2 a) Calculate p12(2) b) Assuming Xo 1 (with probability 1), find the probability that Xn will reach state 2 before it reaches state 4 c) Find msz. d) Is the chain periodic? Irreducible? e) Find the stationary distribution f Approximate the probability that X0 1 g) Find the mean recurrence time for state 1
Answer #1

By using the R-code to find out the stationary distribution is as followsRR Console > tpm-matrix (c (0,1/2, 1/2,0, 1/4, 1/2,0, 1/4, 1/2, o, o, 1/2,0, 1/2, 1/2, o),nrow-4,byrow $ tpm [1, 0.00 0.5 0.5 o.00 [2,1 0.25 0.5 0.0 0.25 [3,1 0.50 0.0 0.0 0.50 14,1 0.00 0.5 0.5 0.00 rowsums (tpm) > a-new (markovchain, transitionMatrix=tpm, states s) > steadyStates (a) I1,1 0.2 0.4 0.2 0.2 >absorbingStates (a) character (0)therefore the stationary distribution is;

(0.2,0.4,0.2,0.2)

Piっ 12 2 Por C2) P1 2 4

b) the answer is 0.4

c) we know that the relation between mean reccurence time and stationary distribution

and which is m32=1/v2

thus, m32=1/0.4=2.5

d) the probability that X1000=1 is 0.2

e) the mean reccurence time for state 1 is 1/v1=5